OpenStudy (anonymous):

Use f(x) = 2x − 5 and g(x) = 3 − x2 to evaluate the expression (f o f)(x) and (g o g)(x)

jimthompson5910 (jim_thompson5910):

f(x) = 2x − 5 f(f(x)) = 2f(x) − 5 f(f(x)) = 2(2x-5) − 5 f(f(x)) = 4x-10 - 5 f(f(x)) = 4x-15